# Discretisation effects of gradient flows in QCD-like theories on the lattice

### 05 Mar 2026 v1

@@ 3 A first look at lattice artefacts
In Fig. 3, we plot the values of the ratio of scales as a function of the lattice spacing (in units of $t_{0}^{\text{Wilson}}$) for the different auxiliary ensembles we generated for Wilson and the DBW2 flow scales. First, we notice that results for different values of $N_{C}$ align nicely with each other at finite lattice spacing, which indicates that the perturbative rescaling performed in Eq. (5) works very well. Second, we observe that the discretisation effects associated with DBW2 flow contribute with opposite sign with respect to the Wilson flow case. Concerning the magnitude of lattice artefacts, although the effect on DBW2 seems somewhat milder, we conclude that they are both of the order of 10%. In the right-hand panel of Fig. (3) we also plot $1-R^{\text{Wilson}}/R^{\text{DBW2}}$, which is expected to vanish in the continuum limit. We attempted a constrained extrapolation to $a=0$ in units of $t_{0}$ through a linear and a quadratic term in $a/\sqrt{t_{0}}$, which confirms the expected behaviour in the continuum limit. A more sophisticated extrapolation is out of the scope of this work, as the coarse values of $a$ combined with the per-cent precision of our scales influence the quality of the fit.
Figure 3: Left-hand side panel: ratio of scales $R=t_{0}/t_{1}$ vs lattice spacing in units of $\sqrt{t_{0}^{\text{Wilson}}}$ for the ensembles in Tab. 1 for different kernel action in the gradient flow discretisation. Right-hand side panel: difference from unity of the ratio $R^{\text{Wilson}}/R^{\text{DBW2}}$ for the same ensembles vs $a/\sqrt{t_{0}}$. Qualitative extrapolation performed as constrained linear + quadratic term in $a$.
We recall that the difference between DBW2 and Wilson lattice artefacts concerns mainly the discretisation effects associated with the smoothening action of the gradient flow [18]. Nevertheless, by only looking at the Wilson flow scales, one can get a rough idea of the order of magnitude of the artefacts we might expect in our future studies. In this regard, we performed the same scale setting analysis on some of the ensembles we plan to use for the spectroscopy study. We selected one ensemble for each $N_{C}$ value by keeping the fake-pion mass [^1] We define the fake pion as the connected piece of the pseudoscalar meson approximately fixed at $\sim 520(20)$ MeV (converting to physical units from the Wilson flow scale using the standard reference value for QCD for comparison).
@@ Acknowledgements
We performed parameter tuning on the Discoverer supercomputer (grant EHPC-REG-2023R01-102) and generated configurations on the LUMI supercomputer (allocation EHPC-EXT-2024E01-038) provided by EuroHPC JU. This work was further supported by a DeiC grant (case 4265-00016A) and the UCloud DeiC Interactive HPC system, managed by the eScience Center at the University of Southern Denmark. S.M. has received funding from the Austrian Science Fund FWF under grant number PAT6443923.

### 26 Mar 2026 v2

@@ 3 A first look at lattice artefacts
In Fig. 3, we plot the values of the ratio of scales as a function of the lattice spacing (in units of $t_{0}^{\text{Wilson}}$) for the different auxiliary ensembles we generated for Wilson and the DBW2 flow scales. First, we notice that results for different values of $N_{C}$ align nicely with each other at finite lattice spacing, which indicates that the perturbative rescaling performed in Eq. (5) works very well. Second, we observe that the discretisation effects associated with DBW2 flow contribute with opposite sign with respect to the Wilson flow case. Concerning the magnitude of lattice artefacts, although the effect on DBW2 seems somewhat milder, we conclude that they are both of the order of 10%. In the right-hand panel of Fig. (3) we also plot $1-R^{\text{Wilson}}/R^{\text{DBW2}}$, which is expected to vanish in the continuum limit. We attempted a constrained extrapolation to $a=0$ in units of $t_{0}$ through a sum of cubic and quadratic terms in $a/\sqrt{t_{0}}$, which confirms the expected behaviour in the continuum limit [18]. A more sophisticated extrapolation is out of the scope of this work, as the coarse values of $a$ combined with the per-cent precision of our scales influence the quality of the fit.
Figure 3: Left-hand side panel: ratio of scales $R=t_{0}/t_{1}$ vs lattice spacing in units of $\sqrt{t_{0}^{\text{Wilson}}}$ for the ensembles in Tab. 1 for different kernel action in the gradient flow discretisation. Right-hand side panel: difference from unity of the ratio $R^{\text{Wilson}}/R^{\text{DBW2}}$ for the same ensembles vs $a/\sqrt{t_{0}}$.
We recall that the difference between DBW2 and Wilson lattice artefacts concerns mainly the discretisation effects associated with the smoothening action of the gradient flow [18]. Nevertheless, by only looking at the Wilson flow scales, one can get a rough idea of the order of magnitude of the artefacts we might expect in our future studies. In this regard, we performed the same scale setting analysis on some of the ensembles we plan to use for the spectroscopy study. We selected one ensemble for each $N_{C}$ value by keeping the fake-pion mass [^1] We define the fake pion through the connected piece of the pseudoscalar correlator. approximately fixed at $\sim 520(20)$ MeV (converting to physical units from the Wilson flow scale using the standard reference value for QCD for comparison).
@@ Acknowledgements
The authors thank Nikolai Husung for useful comments. We performed parameter tuning on the Discoverer supercomputer (grant EHPC-REG-2023R01-102) and generated configurations on the LUMI supercomputer (allocation EHPC-EXT-2024E01-038) provided by EuroHPC JU. This work was further supported by a DeiC grant (case 4265-00016A) and the UCloud DeiC Interactive HPC system, managed by the eScience Center at the University of Southern Denmark. S.M. has received funding from the Austrian Science Fund FWF under grant number PAT6443923.